OpenStudy (anonymous):

A system of linear equations is shown below. 2y = -3x + 4 -2y = 8 + 6x Which of the following describes the graph of this system of linear equations in the standard (x, y) coordinate plane?

OpenStudy (anonymous):

(-4,8) 2y = -3x + 4 -2y = 8 + 6x -------------add 0 = 3x +12 3x = -12 x = -4 sub in one equation 2y = -3x + 4 2y = -3(-4) + 4 2y = 12 + 4 2y = 16 y = 8

OpenStudy (anonymous):

Thank you!!!!